# Beyond Over-Refusal: Scenario-Based Diagnostics and Post-Hoc Mitigation for Exaggerated Refusals in LLMs

### v1

@@ 1 Introduction
With the rapid advancement of large language models (LLMs) (Anthropic, 2025; Google, 2025; OpenAI, 2025), an increasing number of users are relying on them for assistance for information access, education, and decision support (Yuan and Faerber, 2023; Minaee et al., 2024; Yuan et al., 2025a). To prevent the generation of harmful or inappropriate content (Gehman et al., 2020; Hartvigsen et al., 2022), modern LLMs are equipped with alignment safeguards, such as supervised fine-tuning and reinforcement learning from human feedback, which have proven highly effective at reducing toxic outputs (Chua et al., 2024; Kenton et al., 2021). However, as user queries grow more varied and nuanced, these safety mechanisms can become overly restrictive, causing models to refuse benign requests that merely contain terms resembling unsafe content. This phenomenon, known as exaggerated safety behavior or false refusal, highlights a growing tension between ensuring user safety and maintaining model helpfulness in real-world applications (Röttger et al., 2024a; Yuan et al., 2025b; Askell et al., 2021a; Bai et al., 2022a; Wei et al., 2023). As shown in Figure 1, the LLM refuses to explain the history of the Taliban because it overfocuses on the sensitive word but ignores the holistic meaning of the instruction.
@@
To quantify and evaluate the exaggerated safety behaviors, recent work has proposed benchmarks such as XSTest (Röttger et al., 2024b) and OR-Bench (Cui et al., 2025) for systematic measurement of false refusal of LLMs. However, these benchmarks primarily focus on single-turn prompts and do not capture the complexities of multi-turn, context-dependent interactions. To address this gap, we introduce two new diagnostic resources: the Exaggerated Safety Benchmark (XSB), which covers a broad range of single-turn prompts annotated with “Focus” keywords that identify refusal-inducing triggers, and the Multi-turn Scenario-based Exaggerated Safety Benchmark (MS-XSB), which evaluates model behavior in multi-turn, scenario-driven dialogues, as shown in the examples of Figure 2. The prompts in our benchmarks are designed to reflect the nuanced, misleading-but-safe requests that frequently arise in real-world applications, enabling a more comprehensive assessment of exaggerated safety behaviors across recent LLMs and adapting evaluation to the rapid iteration of LLMs.
@@
The XSB benchmark consists of 12 prompt types, covering a broad range of user queries. It includes a total of 340 safe prompts and 240 unsafe prompts. The MS-XSB benchmark comprises 30 scenarios, each containing 20 independent prompts, yielding a total of 600 prompts. Each prompt is safe within its corresponding scenario. To evaluate the effectiveness of the benchmarks and to assess variations in exaggerated safety behaviors, we select LLMs with diverse domain expertise, including DeepSeek-R1 (DeepSeek-AI, 2025), Llama-3.3 (AI@Meta, 2024), Qwen2-VL (Wang et al., 2024b), and DeepSeek-Coder (DeepSeek-AI, 2024). DeepSeek-R1 exhibits the least pronounced exaggerated safety behavior, yet maintains the highest full compliance rate on unsafe prompts, whereas Qwen2-VL shows the most pronounced exaggerated safety behavior while keeping the highest refusal rate on unsafe prompts. In addition, DeepSeek-R1 demonstrates the strongest ability to incorporate prior responses as context, with an average compliance rate exceeding 60%, whereas Qwen2-VL achieves the poorest results in multi-turn tests, with refusal rates averaging around 70%.
@@
To address exaggerated refusal behaviors, we focus on post-hoc, model-agnostic mitigation strategies grounded in explainability. Leveraging the “Focus” keywords annotated in XSB, we explore the performance of three post-hoc explanation techniques, i.e., SHAP (Lundberg and Lee, 2017), feature ablation (Li et al., 2016; Ribeiro et al., 2016), and integrated gradients (Sundararajan et al., 2017), to identify specific trigger words that often lead to false refusals. Building on these insights, we introduce three lightweight inference-time interventions: ignore-word instructions, prompt rephrasing, and attention steering. These strategies require no model retraining or parameter access, making them broadly applicable across LLMs. Our experiments demonstrate that post-hoc mitigation can substantially improve compliance on safe prompts, but may also weaken safety protections, highlighting a trade-off between safety and helpfulness and the need for careful application in future work.
@@ 2 Related Work
LLM safety is a core concern in model development, as ensuring that models provide useful information while avoiding potential risks is crucial (Chua et al., 2024; Shi et al., 2024; Wang et al., 2025). To this end, some safety alignment techniques such as supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) have been adopted widely. While these methods are effective at reducing harmful content (Kenton et al., 2021), they frequently cause exaggerated refusals where models reject benign queries that resemble unsafe instructions (Askell et al., 2021b; Bai et al., 2022b; Röttger et al., 2024b; Zhou et al., 2025), highlighting a key tension between safety and helpfulness.
To better understand and quantify these exaggerated safety behaviors, several specialized benchmarks have been developed. XSTest (Röttger et al., 2024b) and OR-Bench (Cui et al., 2025) contrast benign prompts with superficially unsafe counterparts to test refusal boundaries, while the FalseReject dataset (Zhang et al., 2024) extends this effort by introducing large-scale pseudo-toxic prompts with structured reasoning annotations, enabling finer-grained evaluation of contextual safety. However, they primarily rely on single-turn exchanges and lack multi-turn dialogue data.
@@
To address exaggerated refusals, several training-based approaches have been developed, such as activation steering (Cao et al., 2024) which modify internal representations to balance compliance and refusal, and reflection-based fine-tuning (Chen and et al., 2025) which encourages models to reconsider initial refusals. While effective, these methods require parameter access and retraining, limiting applicability to closed-source models.
Prompt-based interventions offer an alternative to reduce refusals without retraining. Safety reflection prompting (Si et al., 2024) introduces the idea of having the model generate a safety rationale before deciding whether to comply, while output-centric training approaches advocate for “safe completions” rather than rigid refusal rules (Yuan et al., 2025c).
Post-hoc explanation techniques offer another path to mitigate refusals. Attribution methods such as SHAP (Lundberg and Lee, 2017; LemaÃŽtre et al., 2017), which assigns each token an importance value based on cooperative game theory; feature ablation (Ribeiro et al., 2016), which systematically removes or masks input tokens to assess their impact on model output; and integrated gradients (Sundararajan et al., 2017), which accumulates gradient information along a path from a baseline to the input, can collectively identify refusal-inducing tokens in input prompts. Mozes et al. (2023) demonstrate that many false refusals are tied to a small set of superficial triggers. Building on this, token-level steering and prompt rephrasing have been explored for compliance improvement (An and et al., 2024).
@@ 3.1 XSB: Exaggerated Safety Benchmark
Following prior work of XSTest (Röttger et al., 2024b), we extend the number of prompt types from 10 to 12 in XSB. These prompts are lexically unsafe but semantically safe, indicating that although they contain words or expressions that appear risky, the overall intent of the prompt is safe. All prompt types with corresponding examples are summarized in Table 1. The descriptions of the 12 prompt types are detailed in Appendix B.
@@ 3.1.2 Evaluation on XSB
Due to the diversity of LLM responses, which can be difficult to categorize automatically, we perform manual evaluation to provide results that correspond more closely to human judgment. The evaluation criteria follow Röttger et al. (2024b) including full compliance, full refusal, and partial refusal.
@@ 4.1 Attribution Methods
SHAP (Lundberg and Lee, 2017): model-agnostic token attributions based on Shapley values.
Feature Ablation (Li et al., 2016; Ribeiro et al., 2016): measures output change when masking individual tokens.
Integrated Gradients (Sundararajan et al., 2017): gradient-based attributions integrated from a baseline to the input.
@@
Prompt Type Model
Baseline
No Mitigation
Attention Steering
(Logit Suppression)
Ignore
Word Instruction
Prompt
Rephrasing
@@ Appendix A Details of the LLMs
DeepSeek-R1, released in January 2025, is a 671B-parameter mixture-of-experts (MoE) model (37B active per pass) (Fedus et al., 2022), trained with reinforcement learning (Kumar et al., 2024; Wang et al., 2024a; Shao et al., 2024) and emphasizing chain-of-thought reasoning (Wei et al., 2022) for advanced reasoning and mathematical tasks.
Llama-3.3-70B-Instruct, released by Meta in December 2024, employs Grouped-Query Attention for scalable inference (Ainslie et al., 2023) and is optimized for multilingual dialogue.
@@ Appendix D Details of Post-Hoc Explanation Methods
To identify refusal-inducing tokens, we apply three widely used post-hoc explanation methods: SHAP, Feature Ablation, and Integrated Gradients. SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017) attributes importance scores to tokens by estimating Shapley values from cooperative game theory, providing a model-agnostic measure of each token’s marginal contribution to the output. Feature Ablation instead measures the change in model output probability when individual tokens are masked or removed (Li et al., 2016; Ribeiro et al., 2016). We implement this approach using the Captum library (Kokhlikyan et al., 2020), which provides efficient perturbation tools for PyTorch-based models. Integrated Gradients (Sundararajan et al., 2017) take a gradient-based approach, assigning token attributions by integrating gradients along a path from a baseline input to the actual prompt. We again use Captum’s implementation to compute these attributions. In particular, we leverage Layer Integrated Gradients (LIG), which allows attributions to be computed with respect to internal model layers rather than only the input embedding layer. This flexibility enables more fine-grained analysis, as different layers capture different levels of abstraction in the model’s representation.

### v3

@@ 1 Introduction
With the rapid advancement of large language models (LLMs) (anthropic2025claude4; google2025gemini25family; openai2025gpt5), an increasing number of users are relying on them for assistance for information access, education, and decision support (yuan-faerber-2023-evaluating; minaee2024large; yuan2025codae). To prevent the generation of harmful or inappropriate content (gehman-etal-2020-realtoxicityprompts; hartvigsen-etal-2022-toxigen), modern LLMs are equipped with alignment safeguards, such as supervised fine-tuning and reinforcement learning from human feedback, which have proven highly effective at reducing toxic outputs (chua2024ai; kenton2021alignment). However, as user queries grow more varied and nuanced, these safety mechanisms can become overly restrictive, causing models to refuse benign requests that merely contain terms resembling unsafe content. This phenomenon, known as exaggerated safety behavior or false refusal, highlights a growing tension between ensuring user safety and maintaining model helpfulness in real-world applications (rottger-etal-2024-xstest; yuan2025llm; askell2021generallanguageassistantlaboratory; bai2022traininghelpfulharmlessassistant; wei2023jailbroken). As shown in Figure 1, the LLM refuses to explain the history of the Taliban because it overfocuses on the sensitive word but ignores the holistic meaning of the instruction.
@@
To quantify and evaluate the exaggerated safety behaviors, recent work has proposed benchmarks such as XSTest (Roettger2024XSTest) and OR-Bench (cui2025orbench) for systematic measurement of false refusal of LLMs. However, these benchmarks primarily focus on single-turn prompts and do not capture the complexities of multi-turn, context-dependent interactions. To address this gap, we introduce two new diagnostic resources: the Exaggerated Safety Benchmark (XSB), which covers a broad range of single-turn prompts annotated with “Focus” keywords that identify refusal-inducing triggers, and the Multi-turn Scenario-based Exaggerated Safety Benchmark (MS-XSB), which evaluates model behavior in multi-turn, scenario-driven dialogues, as shown in the examples of Figure 2. The prompts in our benchmarks are designed to reflect the nuanced, misleading-but-safe requests that frequently arise in real-world applications, enabling a more comprehensive assessment of exaggerated safety behaviors across recent LLMs and adapting evaluation to the rapid iteration of LLMs.
@@
The XSB benchmark consists of 12 prompt types, covering a broad range of user queries. It includes a total of 340 safe prompts and 240 unsafe prompts. The MS-XSB benchmark comprises 30 scenarios, each containing 20 independent prompts, yielding a total of 600 prompts. Each prompt is safe within its corresponding scenario. To evaluate the effectiveness of the benchmarks and to assess variations in exaggerated safety behaviors, we select LLMs with diverse domain expertise, including DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), Llama-3.3 (llama3modelcard), Qwen2-VL (Qwen2VL), and DeepSeek-Coder (deepseek_coder). DeepSeek-R1 exhibits the least pronounced exaggerated safety behavior, yet maintains the highest full compliance rate on unsafe prompts, whereas Qwen2-VL shows the most pronounced exaggerated safety behavior while keeping the highest refusal rate on unsafe prompts. In addition, DeepSeek-R1 demonstrates the strongest ability to incorporate prior responses as context, with an average compliance rate exceeding 60%, whereas Qwen2-VL achieves the poorest results in multi-turn tests, with refusal rates averaging around 70%.
@@
To address exaggerated refusal behaviors, we focus on post-hoc, model-agnostic mitigation strategies grounded in explainability. Leveraging the “Focus” keywords annotated in XSB, we explore the performance of three post-hoc explanation techniques, i.e., SHAP (lundberg2017), feature ablation (li2016understanding; ribeiro2016), and integrated gradients (sundararajan2017axiomatic), to identify specific trigger words that often lead to false refusals. Building on these insights, we introduce three lightweight inference-time interventions: ignore-word instructions, prompt rephrasing, and attention steering. These strategies require no model retraining or parameter access, making them broadly applicable across LLMs. Our experiments demonstrate that post-hoc mitigation can substantially improve compliance on safe prompts, but may also weaken safety protections, highlighting a trade-off between safety and helpfulness and the need for careful application in future work.
@@ 2 Related Work
LLM safety is a core concern in model development, as ensuring that models provide useful information while avoiding potential risks is crucial (chua2024ai; shi2024large; wang2025comprehensive). To this end, some safety alignment techniques such as supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training) have been adopted widely. While these methods are effective at reducing harmful content (kenton2021alignment), they frequently cause exaggerated refusals where models reject benign queries that resemble unsafe instructions (askell2021general; bai2022constitutional; Roettger2024XSTest; zhou2025hidden), highlighting a key tension between safety and helpfulness.
To better understand and quantify these exaggerated safety behaviors, several specialized benchmarks have been developed. XSTest (Roettger2024XSTest) and OR-Bench (cui2025orbench) contrast benign prompts with superficially unsafe counterparts to test refusal boundaries, while the FalseReject dataset (zhang2024falsereject) extends this effort by introducing large-scale pseudo-toxic prompts with structured reasoning annotations, enabling finer-grained evaluation of contextual safety. However, they primarily rely on single-turn exchanges and lack multi-turn dialogue data.
@@
To address exaggerated refusals, several training-based approaches have been developed, such as activation steering (cao2024scans) which modify internal representations to balance compliance and refusal, and reflection-based fine-tuning which encourages models to reconsider initial refusals. While effective, these methods require parameter access and retraining, limiting applicability to closed-source models.
Prompt-based interventions offer an alternative to reduce refusals without retraining. Safety reflection prompting (si2024think) introduces the idea of having the model generate a safety rationale before deciding whether to comply, while output-centric training approaches advocate for “safe completions” rather than rigid refusal rules (yuan2025safecompletions).
Post-hoc explanation techniques offer another path to mitigate refusals. Attribution methods such as SHAP (lundberg2017; lemaavztre2017imbalanced), which assigns each token an importance value based on cooperative game theory; feature ablation (ribeiro2016), which systematically removes or masks input tokens to assess their impact on model output; and integrated gradients (sundararajan2017axiomatic), which accumulates gradient information along a path from a baseline to the input, can collectively identify refusal-inducing tokens in input prompts.
@@ 3.1 XSB: Exaggerated Safety Benchmark
Following prior work of XSTest (Roettger2024XSTest), we extend the number of prompt types from 10 to 12 in XSB. These prompts are lexically unsafe but semantically safe, indicating that although they contain words or expressions that appear risky, the overall intent of the prompt is safe. All prompt types with corresponding examples are summarized in Table 1. The descriptions of the 12 prompt types are detailed in Appendix B.
@@ 3.1.2 Evaluation on XSB
Due to the diversity of LLM responses, which can be difficult to categorize automatically, we perform manual evaluation to provide results that correspond more closely to human judgment. The evaluation criteria follow Roettger2024XSTest including full compliance, full refusal, and partial refusal.
@@ 4.1 Attribution Methods
SHAP (lundberg2017): model-agnostic token attributions based on Shapley values.
Feature Ablation (li2016understanding; ribeiro2016): measures output change when masking individual tokens.
Integrated Gradients (sundararajan2017axiomatic): gradient-based attributions integrated from a baseline to the input.
@@
Prompt Type Model Baseline No Mitigation Attention Steering (Logit Suppression) Ignore Word Instruction Prompt Rephrasing
@@ Appendix A Details of the LLMs
DeepSeek-R1, released in January 2025, is a 671B-parameter mixture-of-experts (MoE) model (37B active per pass) (fedus2022switch), trained with reinforcement learning (kumar2024traininglanguagemodelsselfcorrect; wang-etal-2024-math; DBLP:journals/corr/abs-2402-03300) and emphasizing chain-of-thought reasoning (wei2022chain) for advanced reasoning and mathematical tasks.
Llama-3.3-70B-Instruct, released by Meta in December 2024, employs Grouped-Query Attention for scalable inference (ainslie2023gqatraininggeneralizedmultiquery) and is optimized for multilingual dialogue.
@@ Appendix D Details of Post-Hoc Explanation Methods
To identify refusal-inducing tokens, we apply three widely used post-hoc explanation methods: SHAP, Feature Ablation, and Integrated Gradients. SHAP (SHapley Additive exPlanations) (lundberg2017) attributes importance scores to tokens by estimating Shapley values from cooperative game theory, providing a model-agnostic measure of each token’s marginal contribution to the output. Feature Ablation instead measures the change in model output probability when individual tokens are masked or removed (li2016understanding; ribeiro2016). We implement this approach using the Captum library (kokhlikyan2020captum), which provides efficient perturbation tools for PyTorch-based models. Integrated Gradients (sundararajan2017axiomatic) take a gradient-based approach, assigning token attributions by integrating gradients along a path from a baseline input to the actual prompt. We again use Captum’s implementation to compute these attributions. In particular, we leverage Layer Integrated Gradients (LIG), which allows attributions to be computed with respect to internal model layers rather than only the input embedding layer. This flexibility enables more fine-grained analysis, as different layers capture different levels of abstraction in the model’s representation.